Use the information below and the 2-stage Growth Model to find Chase’s intrinsic value. Forecast its cash flows from 2019 through 2021. If its current price is $28.52, is it over or under valued?

N

N+1

N+2

N+3

N+4

Beta

1.20

ROE

0.10

0.10

0.10

DPO

0.25

0.45

0.45

Div

0.54
Answer #1

N+1 N+2 N+3 N+4 ROE DPO Dividend growth rate Dividend 0.25 1.0251.045 1.045 0.54 0.55350.578408 0.604436 Calculate the Dividend growth rate (N+2 Stableperiodpayoutratio 1-g/RO DividendPayout Rati o (DPO)-1- 0.25-1-0.10 010x 0.25-1-g 0.025-1-g g = 1+0.025 1.025 Calculate the Dividend growth rate N+3 0.45 1- 0.10 0.10x0.45-1-g 0.045-1-g g 1.045

N+2 2019 N+3 2020 2021 ROE DPO Dividend growth rate Dividend (Cash Flows) Share Price 0.25 1.025 1.045 1.045 $0.54 $0.5535 $0.578408 $0.604436 S7.38S10.52 S11.00 Calculate the stock price (or) Intrinsic value: Di Stockvalue(R)= R -g - $0.5535 0.10-0.025 -S7.38 Di Stockvalue(R)=-g S0.578408 0.10-0.045 -S10.52 Di R -g $0.604436 0.10-0.045 Stock value (P)_ -S10.99 (or)S11.00

S0.60443601.045) S0.5535 S0.578408 S0.604436 0.10-0.045) 1+0.101 1+0.10 2 1+0.10 3 Intrins ic value of stock+ 1+0.10 3 $0.503 18 18+$0.4780231+ S0.454 12 17+$8.628313 -$10.06 The stocks current price is $28.52 and the stocks intrinsic value we got is S10.06. Hence, the current stock price is over valued.

> calculations are wrong

Ashwini Jagtap Sat, Feb 5, 2022 5:09 AM